Solve the formula for h: V=1/3 Bh
Thanks,
Pat

Found 2 solutions by scott8148, checkley77:
Answer by scott8148(6628)

multiplying by 3 __ 3V=Bh __ dividing by B __ 3V/B=h

Answer by checkley77(12844)

V=1/3 Bh
V=BH/3 MULTIPLY BOTH SIDES BY 3.
BH=3V DIVIDE BOTH SIDES BY B.
H=3V/B ANSWER.